OpenStudy (anonymous):

You find that a small business loan in the amount of 50,000 is the amount you need to purchase the restaurant location. After researching banks to find the best interest rate, you find that banks for small businesses offer the best interest rate of 9% interest that compounds monthly for 7 years. 1.What is the monthly payment for this loan? 2.Show the formula that you used and the values used for each variable to calculate the monthly payment. 3.What is the unpaid balance of the loan at the end of the 1st year? 4.Show the formula that you used and the values used for each variable to ca

OpenStudy (anonymous):

rest of the question is in the attachment

OpenStudy (anonymous):

50k= pm/[(1+im)^(1/12)+........................(1+im/12)^(84/12)]

OpenStudy (anonymous):

50000=pm(1+i)^-1+pm(1+i)^-2+.....pm(1+i)^7 let a =1/1+i pma+pma^2+ .................pma^7 pm * {a^1+a^2...........a^7] pm * a(1+a+a^2+.........a^6] pm * a(1-a)/(1-a) (1+a^1+a^2................) pm * a/1-a * {1-a+a-a^2....................a^7 pm * a/1-a * [1-a^7) 50000=PM 1/1-a*(1-a^n 50000 = pm 1/1-a*(1-a^n ) let a=(1/1+im) in your case n = 84 and im assuming you get the 50k and pay it back at the end of the first month